OpenStudy (anonymous):

elimination method x+24=14 x+3=11

OpenStudy (anonymous):

Are you missing the y's?

OpenStudy (anonymous):

no its suppose to be elimination method

OpenStudy (anonymous):

oh wait its wrong hold on x-2y=14 x+3y=11

OpenStudy (anonymous):

x = 64/5 and y = -3/5

OpenStudy (anonymous):

o.k. given your other equations... eq1-eq2 => -5y = 5 => y =-1 =>x =12

OpenStudy (anonymous):

Multiply the bottom equation by -1. This gives you -x -3 = -11, add this to the top equation. This eliminates the x's. Solve for y.

OpenStudy (anonymous):

Once you have solved for y, plug in the y value into either original equation and solve for x.

jhonyy9 (jhonyy9):

x-2y=14 x+3y=11 - subtract --------- 0 -5y=3 --- y=-3/5 than x=14+2(-3/5)=14-(6/5) =(70-6)/5 =64/5 x=64/5 y=-3/5